How many grams of carbon dioxide can form when a mixture of 3.20 g ethylene (C2H4) and 5.90 g of oxygen is ignited, assuming complete combustion to form carbon dioxide and water?

Give the answer again with the correct number of significant figures.
Answer #1

C2H4 (g) + 3O2(g) ------------------> 2CO2(g) + 2H2O(g)

no of moles of C2H4   = W/G.M.Wt

                                     = 3.2/28

                                      = 0.1143 moles

no of moles of O2         = W/G.M.Wt

                                      = 5.9/32

                                       = 0.184 moles

C2H4 (g) + 3O2(g) ------------------> 2CO2(g) + 2H2O(g)

1 moles of C2H4 react with 3 moles of O2

0.1143 moles of C2H4 react with = 3*0.1143/1 = 0.3429 moles of O2 is required

O2 is limiting reactant

3 moles of O2 react with excess of C2H4 to gives 2 moles of CO2

0.184 moles of O2 react with excess of C2H4 to gives = 2*0.184/3 = 0.1226 moles of CO2

mass of CO2 = no of moles * gram molar mass

                       = 0.1226*44   = 5.4g of CO2 >>>>>answer